The breaking strength (in pounds) of a certain new synthetic is normally distributed, with a mean of 175 and a variance of 16. The material is considered defective if the breaking strength is less than 165.8 pounds. What is the probability that a single, randomly selected piece of material will be defective?
Answer by stanbon(75887)

The breaking strength (in pounds) of a certain new synthetic is normally distributed, with a mean of 175 and a variance of 16. The material is considered defective if the breaking strength is less than 165.8 pounds. What is the probability that a single, randomly selected piece of material will be defective?
-----
z(165.8) = (165.8-175)/16 = -0.5750
-----
P(x < 165.8) = P(z < -0.575 = normalcdf(-100,-0.575) = 0.2826
==========
Cheers,
Stan H.